Calculate the pH of a 0.410 M aqueous solution of hypochlorous acid (Hcio, Ka -3.5x10). pH =
Answer #1

Lets write the dissociation equation of HClO

HClO -----> H+ + ClO-

0.41 0 0

0.41-x x x

Ka = [H+][ClO-]/[HClO]

Ka = x*x/(c-x)

Assuming small x approximation, that is lets assume that x can be ignored as compared to c

So, above expression becomes

Ka = x*x/(c)

so, x = sqrt (Ka*c)

x = sqrt ((3.5*10^-8)*0.41) = 1.198*10^-4

since c is much greater than x, our assumption is correct

so, x = 1.198*10^-4 M

so.[H+] = x = 1.198*10^-4 M

we have below equation to be used:

pH = -log [H+]

= -log (1.198*10^-4)

= 3.92

Answer: 3.92

Feel free to comment below if you have any doubts or if this answer do not work